HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

CURRENT PROGRESS IN VARIATIONAL CALCULATIONS OF POLYATOMIC HYDROCARBONS (C_{2}H_{2}, CH_{4}, C_{2}H_{4}, C_{3}H_{4}, AND C_{2}H_{6})

ALENA RODINA, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of

Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia; ANDREI V. NIKITIN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia;

ALEXANDER PROTASEVICH, Laboratory of Atmospheric Absorption Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia; OLEG EGOROV, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia; \underline{MICHAËL REY} , \underline{DOMINIKA VIGLASKA} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687 Reims Cedex 2, France; \underline{VLADIMIR TYUTEREV} , Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia

The prediction of the rovibrational energy levels of the acetylene molecule was studied using both the ab initio^{1} and new empirically refined potential energy surfaces (PESs). The refinement of the PES was done by the fit of the available experimental data on the main ^{12}C_{2}H_{2} isotopologue. The quality of the refined PES was analyzed by the variational calculations of the rovibrational energy levels of four isotopologues (^{13}C_{2}H_{2},^{12}C_{2}D_{2},^{13}CH^{12}CH, and H^{12}C^{12}CD) including their isotopic shifts for some lower energy levels with J≥5. Although under the Born-Oppenheimer approximation the isotopic shift should be independent on the PES, we found that the refined PES gave more accurate results than those based on the ab initio one. The accuracy of the isotopic shift calculation was estimated as an order of magnitude better than the accuracy of the energy levels calculation. In the case of CH_{4}, the rovibrational energy levels in the region of the tetradecade polyad were computed with a high accuracy using the ab initio PES without any empirical corrections. It was found that the prediction of the vibrational energy levels up to 17000 cm^{−1} strongly depended on shape of the PES.

A novel methodology^{2} for computing parameters of the effective both Hamiltonian and dipole moment was applied with respect to C_{2}H_{4}. The obtained for each polyad effective parameters will be used to improve the assignment of the C_{2}H_{4} absorption spectra, measured particularly, around 7000 cm^{−1}.

1doi:10.1016/j.jqsrt.2022.108349, A. V. Nikitin, A. E. Protasevich, A. A. Rodina, M. Rey, A. Tajti, V. G. Tyuterev, J. Quant. Spectrosc. Radiat. Transfer., 292, 108349, (2022).

2doi:10.1063/5.0089097, M. Rey, J. Chem. Phys,156, 224103, (2022).

p-number: p109  Submitted on Tue, 13 Jun 2023 17:27:27 +0200

The full-dimensional ab initio PESs for two structural isomers of the hydrocarbon C_{3}H_{4} (D_{2d}-propadiene and C_{3v}-propyne) were developed up to the 5^{th} order inclusively. The application of the rigid (D_{3d}) approximation for the variational calculations of the line list of C_{2}H_{6} was discussed with respect to the new contraction approach implemented in the variational TENSOR code and presented previously ab initio PES and DMS^{3}.

This work was supported by the project of the Russian Scientific Foundation (RSF, No. 22-42-09022) and Agence Nationale de la Recherche (ANR, French National Research Agency, Grant No. 21-CE30-0053-01)

3O. Egorov, M. Rey, A. V. Nikitin, D. Viglaska, 27th Colloquium on High-Resolution Molecular Spectroscopy, HRMS Cologne 2021 A55. https://hrms21.scg.ch.

p-number: p109  Submitted on Tue, 13 Jun 2023 17:27:27 +0200